SELF-BROADENED WATER FROM 5 TO 15 MICRONS

R. A. TOTH and \underline{L.R. BROWN}, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109 USA.

The self-broadened halfwidths and pressure shifts of water have been obtained using the McMath Fourier transform spectrometer at Kitt Peak National Observatory / Solar National Observatory. Over 1200 vibration-rotation transitions involving the pure rotational and ν_{2} bands of the main isotope have been measured at room temperature up to J = 17 and K_{a} = 9. The widths range from 0.52 to 0.1 cm^{−1}/atm and the pressure shifts from +0.05 to –0.05 cm^{−1}/atm. Additional measurements of isotopic transitions reveal no detectable differences for the widths of the three main isotopes of water vapor.